# Towards Imitation Learning in Real World Unstructured Social Mini-Games in Pedestrian Crowds

### v1

@@ V-A Speedway Dataset
We processed the data by removing agents that were either at a standstill or appeared at the edges of the lidar’s field of view. Furthermore, we clipped each trajectory within a uniform spatio-temporal range ($[-20m,+20m]$ in the x axis and $[-10m,15m]$ in the y axis). Each trajectory is finally processed in the form of a $T\times 4k$ array where $T$ corresponds to the number of frames in that interaction and $4k$ corresponds to the dimension of the joint state, that is, $x_{t}=\left[p^{i},v^{i},\phi^{i}\right]^{\top}$ for $i\in[1,k],p^{i}\in\mathbb{R}^{2},v^{i}\in\mathbb{R}^{1},\phi^{i}\in\mathbb{S}^{1}$. The final processed trajectories in the speedway dataset form the demonstration trajectories, dentoed as $\mathcal{D}$. Features are computed based on distance to the goal position, distance to other agents, and control effort. For MaxEnt IRL and multi-agent MaxEnt IRL, since we do not know the true cost functions of the pedestrians, we directly minimize the root mean square error between the demonstration trajectories and the trajectories generated via the policy $\mathcal{F}^{i}_{\theta^{i}}$.
V-B Baselines
BC-Transformer [46]–In the BC-Transformer baseline, we leverage the power of self-attention mechanisms. The state input consists of a 7-dimensional feature vector, similar to BC-LSTM. The transformer architecture comprises two encoder and decoder layers, each with $16$ and $2$ attention heads. To prevent overfitting, we apply a dropout rate of $0.2$. Training utilizes the AdamW optimizer, with the loss function being Mean Squared Error (MSE). We set the learning rate to $3\times 10^{-4}$ and perform mini-batch training with a batch size of $64$. A $60-40$ cross-validation split is used for evaluation.
PECNet [42]–PECNet jointly predicts the trajectories of all humans in a scene, while respecting social norms. An MLP network is first used to learn to predict sub-goal endpoints for each trajectory. This component encodes past trajectories and future endpoints to sample possible latent future endpoints. Next, the approach samples estimate of the endpoints from the previous step, coupled with a prediction network, to plan the future path using social pooling. This method was selected because of its state of the art performance on the ETH and UCY pedestrian motion planning benchmarks.
V-C Benchmarking IL Methods on the Speedway Dataset
In this evaluation section, we present a comparison of BC, IRL, generative models, and a hybrid IRL-BC approach using the Speedway dataset. The evaluation focuses on their ability to learn policies for planning motion in complex scenarios involving tightly coupled human interactions. Figure 4 visually illustrates the comparisons, where solid lines represent demonstration trajectories and dashed lines represent trajectories generated from the learned cost functions, with arrows indicating the direction of movement. For clarity, we only show the results on the W-E-S dataset and defer the remaining results to the Appendix.
Multi-agent nature of IRL. Our findings demonstrate that multi-agent MaxEnt IRL outperforms MaxEnt IRL (Figures 4(a) and 4(b)), particularly when dealing with diverse trajectories from multiple agents. We further find that while multi-agent MaxEnt IRL outperforms MaxEnt IRL and BC in terms of matching the ground truth demonstrations (Figure 4), it still remains imperfect in capturing diverse behavior modes due to the constant identity covariance as discussed in Section III. Multi-agent MaxEnt IRL, in fact, does manage to capture the different behavior modes perfectly in simulation, as shown in Figure 6(a), when the cost function takes a quadratic form. In the case of MaxEnt IRL, the collective cost function for three agents leads to suboptimal motion planning for the green agent, as observed by the green dashed line in Figure 4(b). Moreover, MaxEnt IRL struggles to capture the motion of the westbound agent (red dashed line) in Figure 4(b). MaxEnt IRL matches the ground truth demonstrations for $1$ agent on average indicating that it will not scale to denser crowds.
BC, hybrid methods. Our findings reveal that Behavior Cloning (BC) and the hybrid BC-IRL approach generally encounter substantial difficulties, marked by their inability to precisely encapsulate the behaviors of all three investigated agents. Particularly, the predictions generated by MaxEnt IRL reside relatively closer to the distribution of the ground truth demonstrations for specific agents, contrasting starkly with predictions from other examined methodologies, which exhibit considerable deviations–spanning several meters–from the groundtruth trajectories.
However, there exists a notable exception when the models are constrained to predict solely the velocity, provided that either the heading angle or an endpoint is supplemented as ground truth. As illustrated in Figure 4(c), BC demonstrates commendable performance, even in a low data environment, when predicting only the velocity. Similarly, PECNet utilizes endpoint conditioning to forecast velocities exclusively. This revelation implies that the limitation of data does not categorically inhibit the feasibility of imitation learning for trajectory planning; rather, conditioning the learning task on supplementary state variables can serve as a valuable approach in such tightly coupled human interactions. These insights underscore the potential of employing strategically conditioned learning models as a viable alternative to navigate the challenges posed by scarce data availability, illuminating the prospects of more nuanced and adaptable learning methodologies in data-limited scenarios.
In Figure 5(a), we quantitatively verify the qualitative results in the previous section. By measuring the root mean square error (RMSE) as a function of time, we observe that MAIRL incurs the lowest RMSE compared to BC, IRL, and PECNet, which implies that MAIRL most accurately recovers the groud truth reward functions of multiple agents, similar to Figure 4(a). In Figure 5(b), we measure the percentage of trajectories below a certain RMSE threshold; steeper curves indicate a more effective learner.
(a) Average RMSE per timestep.
(b) Cumulative RMSE distribution.
Figure 5: Quantitatively benchmarking representative IL approaches: behavior cloning (BC [46]), inverse reinforcement learning (IRL [16] and MAIRL [40]), and generative methods (PECNet [42]). (left) Root mean square error per timestep; larger slope indicates poor long-term planning. (right) Cumulative RMSE distribution demonstrates the percentage of trajectories below a certain RMSE threshold; steeper curves indicate a more effective learner.
Simulation and Partial Conditioning. We hypothesize that the performance of imitation learning (IL) algorithms improves with access to global information. To validate this hypothesis, we conducted experiments running IL in both simulation and with partial conditioning (supplying training with partial state information). As shown in Figures 6(a) and 6(b), the empirical observations suggest a direct correlation between the richness of information and the accuracy and efficacy of various IL algorithms. The results indicate that providing additional state information during training leads to improved performance, validating our hypothesis and highlighting the importance of comprehensive state information in learning algorithms.
In summary, our evaluation reveals the following insights:
  • Multi-agent cost function learning is important. Multi-agent MaxEnt IRL outperforms BC, MaxEnt IRL, generative models, and hybrid IRL-BC in terms of effectively learning policies for multiple agents with diverse behaviors.
  • Conditioning on supplementary state variables or providing global information in simulation leads to better imitation learning.
(a) Multi agent IRL
(b) Behavior cloning
Figure 6: Simulation and Partial conditioning: We show that IL algorithms succeed when provided global information either in simulation (left) or via partial conditioning (right).
VI Conclusion
This research investigated the effectiveness of Imitation Learning (IL) techniques using a novel pedestrian dataset from the University of Texas at Austin campus. The key findings are as follows: $(i)$ Multi-agent MaxEnt IRL outperformed BC, MaxEnt IRL, generative models, and hybrid IRL-BC in learning effective policies for multiple agents with diverse behaviors, highlighting the importance of multi-agent cost function learning in motion planning tasks in such scenarios. $(ii)$ It was also observed that conditioning IL models on additional state variables improved their performance in low data regimes.
VII Future Directions
The purpose of this study was to shed light on IL methods in real world social interaction scenarios. In light of the lessons learned through this work, several lines of future work emerge. First, given the observed success of multi-agent MaxEnt IRL in representing diverse behaviors in SMGs, future work should focus on developing strategies that model the multi-agent interactions, especially addressing challenges such as local observability. The promise shown by conditioning models on additional state variables suggests that investigating advanced conditioning strategies and integrating more sophisticated variables can significantly refine model predictions and improve outcomes. Real-world application and scalability of these models are also essential; testing and refining them in dynamic, real-world conditions such as malls or transit stations, and developing scalable solutions for denser crowds will be critical in assessing the reliability and applicability of these models. Lastly, incorporating external factors like weather and social contexts in modeling pedestrian behaviors can yield more holistic and precise predictions, providing a more comprehensive view of motion dynamics in various environments.

### v2

@@ V-A Data Pre-processing
Each trajectory is finally processed in the form of a $T\times 4k$ array where $T$ corresponds to the number of frames in that interaction and $4k$ corresponds to the dimension of the joint state, that is, $x_{t}=\left[p^{i},v^{i},\phi^{i}\right]^{\top}$ for $i\in[1,k],p^{i}\in\mathbb{R}^{2},v^{i}\in\mathbb{R}^{1},\phi^{i}\in\mathbb{S}^{1}$. The final processed trajectories in the speedway dataset form the demonstration trajectories, denoted as $\mathcal{\widehat{D}}$. Features are computed based on distance to the goal position, distance to other agents, and control effort.
V-B Baselines and Evaluation Metric
We compare our approach, which we call MAIRL (multi-agent IRL), with the following baselines [^3] Note that these baselines were selected by identifying state-of-the-art methods that $(i)$ were compatible with our data format and $(ii)$ had open source implementations.: single-agent MaxEnt IRL [26], a state-of-the-art behavior cloning (trajectory prediction) algorithm [28], and a generative modeling algorithm (PECNet [30]). We directly use the opensource software for all these baselines and train them on our dataset to present a direct evaluation. We set the learning rate to $3\times 10^{-4}$ and perform mini-batch training with a batch size of $64$. A $60-40$ cross-validation split is used for evaluation. To prevent overfitting, we apply a dropout rate of $0.2$. Training utilizes the AdamW optimizer with the Mean Squared Error (MSE) loss function. All training was performed on an NVIDIA GeForce RTX 2080 GPU machine and the average time to train the baselines was in the order of minutes. Since we do not know the true cost functions of the pedestrians, we are unable to measure some important social navigation metrics such as “comfort”, although we acknowledge those are important as well, and we leave that for future work. Instead, we directly minimize the root mean square error between the demonstration trajectories and the trajectories generated via the policy $\mathcal{F}^{i}_{\theta^{i}}$. This is a standard evaluation protocol followed in the navigation literature [26, 29].
V-C Results and Discussion
Method ETH Hotel Univ Zara1 Zara2 Speedway Average
Social-GAN [14] $0.81/1.52$ $0.72/1.61$ $0.60/1.26$ $0.34/0.69$ $0.42/0.84$ $0.58/1.18$
SoPhie [15] $0.70/1.43$ $0.76/1.67$ $0.54/1.24$ $0.30/0.63$ $0.38/0.78$ $0.54/1.15$
CGNS [74] $0.62/1.40$ $0.70/0.93$ $0.48/1.22$ $0.32/0.59$ $0.35/0.71$ $0.49/0.97$
S-BiGAT* [16] $0.69/1.29$ $0.49/1.01$ $0.55/1.32$ $0.30/0.62$ $0.36/0.75$ $0.48/1.00$
MATF [75] $1.01/1.75$ $0.43/0.80$ $0.44/0.91$ $0.26/0.45$ $0.26/0.57$ $0.48/0.90$
Next* [76] $0.73/1.65$ $0.30/0.59$ $0.60/1.27$ $0.38/0.81$ $0.31/0.68$ $0.46/1.00$
SR-LSTM [77] $0.63/1.25$ $0.37/0.74$ $0.51/1.10$ $0.41/0.90$ $0.32/0.70$ $0.45/0.94$
STGAT [78] $0.65/1.12$ $0.35/0.66$ $0.52/1.10$ $0.34/0.69$ $0.29/0.60$ $0.43/0.83$
GOAL-GAN* [79] $0.59/1.18$ $0.19/0.35$ $0.60/1.19$ $0.43/0.87$ $0.32/0.65$ $0.43/0.85$
SGCN [80] $0.63/1.03$ $0.32/0.55$ $0.37/0.70$ $0.29/0.53$ $0.25/0.45$ $0.37/0.65$
MANTRA [81] $0.48/0.88$ $0.17/0.33$ $0.37/0.81$ $0.27/0.58$ $0.30/0.67$ $0.32/0.65$
Transformer [17] $0.61/1.12$ $0.18/0.30$ $0.35/0.65$ $0.22/0.38$ $0.17/0.32$ $0.31/0.55$
SMEMO [18] $0.39/0.59$ $0.14/0.20$ $0.23/0.41$ $0.19/0.32$ $0.15/0.26$ $0.22/0.35$
Introvert [82] $0.42/0.70$ $0.11/0.17$ $0.20/0.32$ $0.16/0.27$ $0.16/0.25$ $0.21/0.34$
PCCSNet [19] $0.28/0.54$ $0.11/0.19$ $0.29/0.60$ $0.21/0.44$ $0.15/0.34$ $0.21/0.42$
MaxEnt IRL [26] $0.76$ / $1.48$ $0.68$ / $1.74$ $0.51$ / $1.21$ $0.37$ / $0.69$ $0.29$ / $0.75$ $0.40/0.80$ $0.50/1.11$
PECNet [83] $0.54/0.87$ $0.18/0.24$ $0.35/0.60$ $0.22/0.39$ $0.17/0.30$ $0.35/0.75$ $0.30/0.51$
SceneTransformer [20] $0.50$ / $0.76$ $0.14$ / $0.20$ $0.29$ / $0.42$ $0.22$ / $0.36$ $0.16$ / $0.27$ $0.33/0.72$ $0.27/0.46$
AgentFormer [21] $0.45/0.75$ $0.14/0.22$ $0.25/0.45$ $0.18/0.30$ $0.14/0.24$ $0.31/0.76$ $0.24/0.44$
HST [22] $0.41$ / $0.73$ $0.10$ / $0.14$ $0.24$ / $0.44$ $0.17$ / $0.30$ $0.14$ / $0.24$ $0.35/0.68$ $0.24/0.42$
LB-EBM [84] $0.30/0.52$ $0.13/0.20$ $0.27/0.52$ $0.20/0.37$ $0.15/0.29$ $0.32/0.81$ $0.23/0.45$
Trajectron++ [23] $0.39/0.83$ $0.12/0.19$ $0.22/0.43$ $0.17/0.32$ $0.12/0.25$ $0.29/0.75$ $0.22/0.47$
Expert-Goals [24] $0.37/0.65$ $0.11/0.15$ $0.20/0.44$ $0.15/0.31$ $0.12/0.25$ $0.29/0.69$ $\mathbf{0.21/0.41}$
MAIRL $0.39$ / $0.71$ $0.13$ / $0.18$ $0.23$ / $0.32$ $0.19$ / $0.29$ $0.15$ / $0.28$ $\mathbf{0.26/0.60}$ $0.23$ / $0.39$
TABLE II: Quantitative results of our proposed method compared to state-of-the-art baselines on five benchmark datasets. All numbers are in meters.
Dataset PecNET [30] SMEMO [18] OpenTraj [85] JRDB_Traj [86] MaxEnt IRL [26] Social-pose [87] HST [22] MAIRL
JRDB [71] $3.939$ $2.671$ $3.498$ $2.646$ $4.249$ $2.558$ $2.506$ $2.640$
SCAND [6] $0.55/1.20$ $0.50/1.10$ $0.44/0.90$ $0.36/0.81$ $0.42/0.85$ $0.23/0.67$
TABLE III: EFE results on the JRDB and SCAND dataset.
V-C1 Results
In Figure 1, we plot the root mean square error (RMSE) between the planned and ground truth trajectories as a function of the timesteps. Here, a timestep refers to the LiDAR’s capture frequency, and not in terms of seconds. We observe that MAIRL incurs the lowest RMSE compared to behavior cloning, MaxEnt IRL, and PECNet, which implies that MAIRL most accurately recovers the ground truth reward functions of multiple agents. In Figure 3, we plot a cumulative distribution frequency (CDF) that measures the percentage of trajectories that fall below a certain RMSE threshold; steeper curves imply a more effective algorithm as it indicates that that the algorithm was able to generate more low-RMSE trajectories. The CDF plot is useful because the RMSE curve alone may not convey the full picture. For example, a baseline could yield a poor RMSE, and be declared as weak algorithm, but a low RMSE could result from one or two outlier trajectories with an particularly high RMSE, thereby skewing the average against its favor.
[…]
V-C2 Discussion
Our findings demonstrate that multi-agent MaxEnt IRL outperforms MaxEnt IRL, particularly when dealing with diverse trajectories from multiple agents. Due to the single-agent nature of MaxEnt IRL, it aligns the predicted trajectories with the ground truth demonstrations for at most $1$ agent (the blue agent), and fails for the green (extreme deviation) and the red (not reached goal) agents (Figure 4(b)). Behavior cloning encounters substantial difficulties, on account of the inherently non-linear pedestrian trajectories. Note that this is in contrast to many successful results in the autonomous driving domain, where agents follow structured motion along straight lanes and roads. Furthermore, our findings suggest there exists a notable exception when the models are constrained to predict solely the velocity, provided that either the heading angle or an endpoint is supplemented as ground truth to the training model (as is in the case of PECNet). We omit the results here due to lack of space, but we observed that behavior cloning performs better, when predicting only the velocity, although we acknowledge that providing the heading angle as ground truth is impractical in real world scenarios. Finally, we note that even though MAIRL outperforms all three baselines in terms of aligning with the ground truth demonstrations, it still remains imperfect in its own alignment due to the reduction in accuracy resulting from the Tractability-Rationality Trade-off, as discussed in Section III.
VI Conclusion and Open Questions
We proposed a novel approach for multi-agent MaxEnt inverse reinforcement learning for real world unstructured pedestrian crowds. Our key insight was a mathematical trick that enabled tractability of multi-agent MaxEnt IRL at the cost of a slight reduction in accuracy. We used a dynamic game theoretic solver to compute optimal policies that maximize the reward functions learned via our approach. We evaluated this new approach using a dataset collected at a busy intersection on a University campus and showed that despite the accuracy trade-off, our approach outperforms state-of-the-art imitation learning baselines such as single-agent MaxEnt IRL, behavior cloning, and generative modeling. Our findings highlighted two key insights: $(i)$ our approach yields the lowest error between the policy-generated trajectories and the ground truth trajectories, $(ii)$ our approach has the highest percentage of low-RMSE trajectories. Future work includes finding ways to solve Equation (17) in unstructured pedestrian crowds without relaxing bounded rationality. Another open problem is to compute an exact bound for the tractability-rationality tradeoff trick (Section IV). In addition, incorporating external factors like social contexts and the environment in modeling pedestrian behaviors can yield more holistic and precise predictions, providing a more comprehensive view of motion dynamics in various environments.